Problem 2: We consider an auction with two bidders. Each bidder's valuation v is drawn as follows: 

• With probability the valuation v is distributed uniformly between 0 and 2.

 • With probability 3 the valuation v is distributed uniformly between a and 1. 

1. If we run a first price auction, what is the optimal bid function of a bidder with valuation v?

 2. What is the expected revenue of the seller? 

3. Suppose now that there is only one bidder and that the seller runs a second price auction with a reserve price. Calculate the optimal reserve price.


1
Expert's answer
2021-02-09T18:18:15-0500
"Solution"


1.Let "B_i" be agent i's bid and "\pi's" agent i's profit. If "B_i\geq\ v_i,\ then\ \pi\leq0,\"then assuming rationality "B_i\leq V_i"

Thus "\pi=0"

If you want to maximize your expected profit (hence your valuation of money is 

risk-neutral), then your maximum bid is "V-2B=0\implies\ B=\frac{v}{2}" where "B" is the bid.


2. Here there are two options in place. One, a second-price auction, the seller is committed to collecting less money, because of charging the second-highest bid. On the other hand, in a first price auction, the bidders reduce their bids, in return reducing what the seller can collect.

Therefore, the seller's expected revenue is

"=(\frac{n-2}{n})(\frac{n}{n+2})=\frac{n-2}{n+2}\\
Where\ n=2\\
=\frac{2-2}{2+2}=\frac{0}{4}\\
=0"

3.Due to the reserve price the item will be on the highest bidder, only if the highest bid is above "r", otherwise the item should not be sold. In the second price auction, if there is a winner, he pays the maximum of the second-place bid and the reserve price "r".

If the item is worth "v" to the seller, then r=v with a probability of "2-r", the bidders' value is above "r" and this item will be sold at a price "r". Given a probability "r", the bidders value will be below "r", and the seller will hold the item with a payoff of "v=0". Therefore, the seller expects revenue which is "r(2-r)," this is maximized at "r=\frac{1}{2}"